OpenStudy (anonymous):

How do I evaluate tan(17pi/12) using the sum identity of tangent?

OpenStudy (anonymous):

17pi / 12 = 2pi / 3 + 3pi / 4

OpenStudy (anonymous):

I have that part already I just don't know how to apply that to the sum identity

OpenStudy (anonymous):

tan (a + b) = (tan a + tan b) / (1 - tan a tan b)

OpenStudy (anonymous):

In your question, a = 2pi / 3 and b = 3pi / 4

OpenStudy (anonymous):

would it be \[(\tan(\frac{ 2\pi}{ 3 })+\tan(\frac{ 3\pi }{ 4 })/(1-(\tan(\frac{ 2\pi}{ 3 })\tan(\frac{ 3\pi }{ 4 })\]

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

so would I solve that by calculator and that would be my answer?

OpenStudy (anonymous):

Solve it part by part

OpenStudy (anonymous):

tan (2pi / 3) = - sqrt 3 tan (3pi / 4) = -1

OpenStudy (anonymous):

Substitute those values and simplify your fraction

OpenStudy (anonymous):

okay thank you! :)